# Hydrogen Diffusion in Magnesium Using Machine Learning Potentials

### v1

@@ 3 Results and Discussions
Figure 5: Diffusion path of a representative hydrogen atom during 100 ps at 673K, within the magnesium bulk, depicted using a color-gradient line to represent the progression of time, from blue to red. The black circles highlight the interstitial regions where H atoms tend to oscillate around Mg lattice site before continuing their diffusing trajectory.
Figure 6: Number of hydrogen pairs below a threshold distance of 3 Å  averaged over 1 ps (in red), and the average value over the entire simulation (black dashed line).
The third significant peak in the RDF is observed around 3 Å in the Mg-Mg curve, consistently with the typical magnesium distances in hcp structures. Multiple smaller peaks indicate further neighbor interactions in the crystal lattice. Analogous results were found in the RDF at 300 K and 480 K, where the peaks are sharper due to the reduced effect of thermal motion, see Supplementary Figure 3. In particular, higher pronounced RDF at lower temperature indicates that hydrogen tend to spend more time in the vicinity of Magnesium, and diffuse less in the crystal structure.
@@ 4 Conclusion
Our investigation enabled to thoroughly characterize the kinetic properties and mobility of hydrogen within a structured environment, such as pure magnesium, across various temperatures, through a rigorous and efficient methodology. We performed a systematic and comparative study of ML-based interatomic potential MD schemes, particularly focusing on Bayesian versus equivariant and graph neural networks (MACE and CHGNet), under different training modes (universal and fine-tuned). The results were validated by estimates of the diffusion coefficient values, which showed excellent agreement with experimental data. The obtained results proved that the VASP-DFT configurations, collected during the MLFF on-the-fly training, represent a complete set for the accurate modeling of inter-atomic interactions, between hydrogen and magnesium atoms. This strategy could be consistently applied to generate a comprehensive dataset for the proper training of existing or forthcoming potentials. In fact, we identified the limitations of pretrained UIPs in studying materials with diffusing defects, due to the absence of representative high temperatures, defective and metastable states in their datasets. Particularly, we demonstrated the ability of state-of-the-art machine learning models to achieve DFT-level accuracy, after fine-tuning on actively learned DFT-configurations, highlighting the importance of focusing on efficient dataset-building methods and their quality. Specifically, the importance of including defected and transition-state configurations, as well as the role of transfer-learning, allowing pre-trained solutions to adapt to new systems. These progresses not only improve our understanding of hydrogen interactions in magnesium, but also pave the way for future research into a broader range of multi-component systems and defected compositions. This is especially significant for systems with complex potential energy surfaces, where traditional ab-initio methods become impractical. Successfully modeling the hydrogen diffusion mechanism in magnesium via machine learning-accelerated molecular dynamics could facilitate the study and discovery of new, more efficient materials for hydrogen storage and beyond it, contributing to the transition towards a greener and more sustainable energy future.
@@ Acknowledgments
C. F. and A. A. acknowledges the ”Doctoral College Advanced Functional Materials – Hierarchical Design of Hybrid Systems DOC 85 doc.funds” funded by the Austrian Science Fund (FWF) and by the Vienna Doctoral School in Physics (VDSP). For Open Access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. D. M. and S. P. were supported by the European Union Horizon 2020 research and innovation program under Grant Agreement No. 857470, from the European Regional Development Fund under the program of the Foundation for Polish Science International Research Agenda PLUS, grant No. MAB PLUS/2018/8, and the initiative of the Ministry of Science and Higher Education ’Support for the activities of Centers of Excellence established in Poland under the Horizon 2020 program’ under agreement No. MEiN/2023/DIR/3795. L.P. and C.F. acknowledges the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.3 - Project NEST (Network 4 Energy Sustainable Transition) of Ministero dell’Università e della Ricerca (MUR), funded by the European Union – NextGenerationEU. L.L. and C.F. acknowledges the NRRP, CN-HPC grant no. (CUP) J33C22001170001, SPOKE 7, of MUR, funded by the European Union – NextGenerationEU. The computational results were obtained using the Vienna Scientific Cluster (VSC) and the LEONARDO cluster through the ISCRA initiative.
@@
Conceptualisation: C.F., A. A., D. M. and L. L. Methodology and calculations: A. A., D. M. and L. L. Supervision: C.F. Writing: A.A , D.M. , L. L. and C. F. Discussion and reviewing: All authors

### v2

@@ 3 Results and Discussions
Figure 5: Diffusion path of a representative hydrogen atom in MgH0.0625 during 100 ps at 673K, depicted using a color-gradient line to represent the progression of time, from blue to red. The black circles highlight the interstitial regions where H atoms tend to oscillate around Mg lattice site before continuing their diffusing trajectory.
The analysis of hydrogen pair ($n_{p}$) distributions presented in Fig. 6-b) provides a clear insight into the dependence of pair formation on temperature and concentration. At elevated temperatures, such as 480 K and 673 K, the simulated distributions align closely with the theoretical Poisson distribution, calculated as:
$f(n_{p};\lambda)=\frac{e^{-\lambda}\lambda^{n_{p}}}{n_{p}!},$
where $\lambda$ represents the average number of pairs. This agreement demonstrates the random nature of hydrogen pairing, indicating a lack of correlation and independence from the initial atomic configuration. Furthermore, the convergence of $\lambda$ values for different concentrations as temperature increases underscores this intrinsic randomness. Conversely, at 300 K, a deviation from the Poissonian behavior emerges, with $\lambda$ showing a stronger dependence on hydrogen concentration. This behavior can be attributed to the reduced mobility of hydrogen atoms at low temperatures, where the thermal energy is insufficient to overcome the energy barriers for interstitial site transitions. Consequently, the hydrogen atoms remain in proximity to their initial positions, leading to non-equilibrated pair distributions within the simulation timescale. These results suggest that, while the system approaches a random distribution at higher temperatures, longer simulations would be required to achieve equilibration and Poissonian behavior at room temperature for low-mobility regimes.
Figure 6: Results obtained from runs at different hydrogen content using MACE_FT. a) Diffusion coefficients at different temperature and H concentration, on top, and obtained activation energy at different concentration, on the bottom. The target experimental value from [19] is reported as the dashed line in every plot. b) Comparison of the simulated hydrogen pair distributions ($n_{p}$ occurrences, bars) with theoretical Poisson distributions (points and dashed lines) at three temperatures (300 K, 480 K, and 673 K) for varying hydrogen concentrations. The lower panel shows the $\lambda$ parameter (average number of pairs) as a function of temperature and concentration.
@@ 4 Conclusion
Our investigation enabled to thoroughly characterize the kinetic properties and mobility of hydrogen within a structured environment, such as pure magnesium, across various temperatures through a rigorous and efficient methodology. We performed a systematic and comparative study of ML-based interatomic potential MD schemes, particularly focusing on Bayesian versus equivariant and graph neural networks (MACE and CHGNet), under different training modes (universal and fine-tuned). The results were validated by estimates of the diffusion coefficient and the activation energy values, which showed excellent agreement with experimental data. The obtained results proved that the VASP-DFT configurations, collected during the MLFF on-the-fly training, represent a complete set for the accurate modeling of inter-atomic interactions, between hydrogen and magnesium atoms. This strategy could be consistently applied to generate a comprehensive dataset for the proper training of existing or forthcoming potentials. In fact, we identified the limitations of pretrained UIPs in studying materials with diffusing defects, due to the absence of representative high temperatures, defective and metastable states in their datasets. Particularly, we demonstrated the ability of state-of-the-art machine learning models to achieve DFT-level accuracy, after fine-tuning on actively learned DFT-configurations, highlighting the importance of focusing on efficient dataset-building methods and their quality. Specifically, the importance of including defected and transition-state configurations, as well as the role of transfer-learning, allowing pre-trained solutions to adapt to new systems. The obtained data offer valuable new insights into the collective dynamical properties at varying hydrogen concentrations, accurately predicting a decrease in hydrogen mobility as concentration increases. Additionally, the statistical analysis of hydrogen pair distributions indicates the system approaches a random distribution at higher temperatures. However, longer simulations would be necessary to achieve equilibration and Poissonian behavior at room temperature in low-mobility regimes.
These progresses not only improve our understanding of hydrogen interactions in magnesium, but also pave the way for future research into a broader range of multi-component systems and defected compositions. This is especially significant for systems with complex potential energy surfaces, where traditional ab-initio methods become impractical. Successfully modeling the hydrogen diffusion mechanism in magnesium via machine learning-accelerated molecular dynamics could facilitate the study and discovery of new, more efficient materials for hydrogen storage and beyond it, contributing to the transition towards a greener and more sustainable energy future.
@@ Acknowledgments
C. F. and A. A. acknowledges the ”Doctoral College Advanced Functional Materials – Hierarchical Design of Hybrid Systems DOC 85 doc.funds” funded by the Austrian Science Fund (FWF) and by the Vienna Doctoral School in Physics (VDSP). For Open Access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. D. M. and S. P. were supported by the European Union Horizon 2020 research and innovation program under Grant Agreement No. 857470, from the European Regional Development Fund under the program of the Foundation for Polish Science International Research Agenda PLUS, grant No. MAB PLUS/2018/8, and the initiative of the Ministry of Science and Higher Education ’Support for the activities of Centers of Excellence established in Poland under the Horizon 2020 program’ under agreement No. MEiN/2023/DIR/3795. L.P. and C.F. acknowledges the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.3 - Project NEST (Network 4 Energy Sustainable Transition) of Ministero dell’Università e della Ricerca (MUR), funded by the European Union – NextGenerationEU. L.L. and C.F. acknowledges the NRRP, CN-HPC grant no. (CUP) J33C22001170001, SPOKE 7, of MUR, funded by the European Union – NextGenerationEU. The computational results were obtained using the Vienna Scientific Cluster (VSC) and the LEONARDO cluster. We acknowledge access to LEONARDO at CINECA, Italy, via an AURELEO (Austrian Users at LEONARDO supercomputer) project
@@
The training dataset can be openly accessed at https://figshare.com/articles/dataset/_b_ML_AB_b_/27922470?file=50852211.
@@
Conceptualisation: C.F., A. A., D. M. and L. L. Methodology and calculations: A. A., L.L. and D. M. Supervision: C.F. Writing: A.A , D.M. , L. L. and C. F. Discussion and reviewing: All authors